OpenStudy (anonymous):

7 3/4=x/8 A. 1/4 B.62 C. 1/62 D. 4 HElp only this quesion

OpenStudy (alexandervonhumboldt2):

first convert into improper fraction: \[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[7\frac{ 3 }{ 4 }=\frac{ 7*4+3 }{ 4 }=?\]

OpenStudy (anonymous):

b.x=62

OpenStudy (alexandervonhumboldt2):

please explain your answwer, do not give out direct answers

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i was just thinking y would u ask a question if you dont want the answer

OpenStudy (alexandervonhumboldt2):

person asks question to learn, not for an asnwer

OpenStudy (anonymous):

ok so just show them how to get the answer

OpenStudy (anonymous):